# Recent Advances in Speech Language Models: A Survey

### v1

@@ 7 Challenges and Future Directions
While SpeechLMs have demonstrated impressive abilities, the research in this area is still in its infancy. In this section, we survey challenges, unsolved questions, and possible directions for future research in the study of SpeechLMs.
7.1 Understanding Different Component Choices
@@
Current research on SpeechLMs encompasses key components such as speech tokenizers, language models, and vocoders, each offering a diverse range of options. While some studies have compared various component choices—primarily focusing on speech tokenizers—the comparisons tend to be limited in scope and depth Lakhotia et al. (2021); Rubenstein et al. (2023). Consequently, there remains a significant gap in understanding the advantages and disadvantages of different component selections. Therefore, studies aimed at comprehensively comparing these choices are essential. Such an investigation would yield valuable insights and serve as a guide for selecting more efficient components when developing SpeechLMs.
7.2 End-to-End Training
@@
Although SpeechLMs can generate speech directly without relying on text signals, they still need to train the three components separately. This separate optimization may hinder the model’s overall potential. Consequently, it would be worthwhile to investigate whether training can be conducted in an end-to-end manner, allowing gradients to be back-propagated from the vocoder’s output to the tokenizer’s input. By exploring this fully end-to-end approach, we could potentially enable SpeechLMs to produce more coherent, contextually relevant, and high-fidelity speech outputs.
7.3 Real-Time Speech Generation
@@
Enabling real-time speech generation is crucial in SpeechLM as it fosters a more interactive way of engaging with humans. However, the most adopted approaches described in section 3 still result in noticeable delays between input and output speech generation. This delay occurs because a typical vocoder must wait for the entire sequence of output tokens to be generated by the language model before functioning, making it the most time-consuming process in the inference pipeline. One potential solution to improve latency is to develop a streamable vocoder, allowing it to begin synthesizing output speech while the language model generates output speech tokens. Another option could involve the SpeechLM autonomously generating audio samples in waveform. Overall, this area of real-time speech generation remains underexplored and requires further investigation.
7.4 Safety Risks in SpeechLMs
@@
Safety is a highly significant subject in the field of Machine Learning, particularly when it comes to large-scale generative AI models. While there has been extensive research on safety concerns in TextLMs, the safety issues in SpeechLMs have not been thoroughly investigated. The safety challenges in SpeechLMs present both similarities and unique aspects compared to TextLMs, as highlighted in OpenAI’s recent report on the safety issues of GPT-4o’s voice model OpenAI (2024). Therefore, it is crucial for future research to explore safety vulnerabilities in SpeechLMs and develop safer SpeechLMs.
@@
Primary concerns for the safety issues in SpeechLMs include but are not limited to toxicity and privacy. Toxicity refers to the harmful nature of the content generated by SpeechLMs. For instance, these models might produce semantically dangerous content, such as instructions for making explosives. Additionally, they could generate acoustically inappropriate content, like erotic speech OpenAI (2024), which presents a unique challenge. Privacy involves the risk of revealing personal information from the speech input after it has been processed by a SpeechLM. For example, the model might infer the speaker’s identity based on the semantic content or acoustic features of the input. Even more concerning is the potential for the model to make biased inferences about the speaker, such as their ethnicity or religious beliefs, based on insufficient (e.g., acoustic) information OpenAI (2024).
7.5 Performance on Rare Languages
@@
SpeechLMs directly model speech data, which allows them to more effectively handle “low-resource” languages compared to TextLMs. “Low-resource” languages are those that lack extensive textual data, making it challenging for TextLMs to model them efficiently. In contrast, SpeechLM provides a better solution by modeling the speech data of these “low-resource” languages, which often have more available audio data than text Lakhotia et al. (2021). Therefore, future research could focus on training SpeechLMs in “low-resource” languages or dialects to expand their capabilities.
8 Conclusions
@@
This survey provides a comprehensive overview of recent advancements in Speech Language Models (SpeechLMs). We begin by addressing the limitations of the naive framework that combines Automatic Speech Recognition (ASR), Large Language Models (LLMs), and Text-to-Speech (TTS) systems for voice interactions. Next, we highlight the key advantages offered by SpeechLMs. Following this, we explore the architectures of SpeechLMs, detailing the components involved and their training recipes. We also discuss their capabilities in various downstream applications as well as their various evaluation methods. Finally, we identify the major challenges in developing SpeechLMs and outline potential directions for future research. We hope this survey will illuminate the field and assist the research community in creating more powerful Speech Language Models.

### v2

@@ VII Challenges and Future Directions
While SpeechLMs have demonstrated impressive abilities, the research in this area is still under explored. In this section, we survey challenges, unsolved questions, and possible directions for future research in the study of SpeechLMs.
VII-A Understanding Different Component Choices
@@
Current research on SpeechLMs encompasses key components such as speech tokenizers, language models, and vocoders, each offering a diverse range of options. While some studies have compared various component choices—primarily focusing on speech tokenizers—the comparisons tend to be limited in scope and depth [37, 39]. Consequently, there remains a significant gap in understanding the advantages and disadvantages of different component selections. Therefore, studies aimed at comprehensively comparing these choices are essential. Such an investigation would yield valuable insights and serve as a guide for selecting more efficient components when developing SpeechLMs.
VII-B End-to-End Training
@@
Although SpeechLMs can generate speech directly without relying on text signals, some studies train the three components separately. This separate optimization may hinder the model’s overall potential. Consequently, it would be worthwhile to investigate whether training can be conducted in an end-to-end manner, allowing gradients to be back-propagated from the vocoder’s output to the tokenizer’s input. By exploring this fully end-to-end approach, we could potentially enable SpeechLMs to produce more coherent, contextually relevant, and high-fidelity speech outputs.
VII-C Real-Time Speech Generation
@@
Enabling real-time speech generation is crucial in SpeechLM as it fosters a more interactive way of engaging with humans. However, the most adopted approaches described in section III still result in noticeable delays between input and output speech generation. This delay occurs because a typical vocoder must wait for the entire sequence of output tokens to be generated by the language model before functioning, making it the most time-consuming process in the inference pipeline. One potential solution to improve latency is to develop a streamable pipeline, allowing the speech input and output to be processed and generated in chunks. Another option could involve the SpeechLM autonomously generating audio samples in waveform. Overall, this area of real-time speech generation remains underexplored and requires further investigation.
VII-D Safety Risks in SpeechLMs
@@
Safety is a highly significant subject in the field of Machine Learning, particularly when it comes to large-scale generative AI models. While there has been extensive research on safety concerns in TextLMs, the safety issues in SpeechLMs have not been thoroughly investigated. The safety challenges in SpeechLMs present both similarities and unique aspects compared to TextLMs, as highlighted in OpenAI’s recent report on the safety issues of GPT-4o’s voice model [136]. Therefore, it is crucial for future research to explore safety vulnerabilities in SpeechLMs and develop safer SpeechLMs.
@@
Primary concerns for the safety issues in SpeechLMs include but are not limited to toxicity and privacy. Toxicity refers to the harmful nature of the content generated by SpeechLMs. For instance, these models might produce semantically dangerous content, such as instructions for making explosives. Additionally, they could generate acoustically inappropriate content, like erotic speech [136], which presents a unique challenge. Privacy involves the risk of revealing personal information from the speech input after it has been processed by a SpeechLM. For example, the model might infer the speaker’s identity based on the semantic content or acoustic features of the input. Even more concerning is the potential for the model to make biased inferences about the speaker, such as their ethnicity or religious beliefs, based on insufficient (e.g., acoustic) information [136].
VII-E Performance on Rare Languages
@@
SpeechLMs directly model speech data, which allows them to more effectively handle “low-resource” languages compared to TextLMs. “Low-resource” languages are those that lack extensive textual data, making it challenging for TextLMs to model them efficiently. In contrast, SpeechLM provides a better solution by modeling the speech data of these “low-resource” languages, which often have more available audio data than text [37]. Therefore, future research could focus on training SpeechLMs in “low-resource” languages or dialects to expand their capabilities.
VIII Conclusions
@@
This survey provides a comprehensive overview of recent advancements in Speech Language Models (SpeechLMs). We begin by addressing the limitations of the naive framework that combines Automatic Speech Recognition (ASR), Large Language Models (LLMs), and Text-to-Speech (TTS) systems for voice interactions. Next, we highlight the key advantages offered by SpeechLMs. Following this, we explore the architectures of SpeechLMs, detailing the components involved and their training recipes. We also discuss their capabilities in various downstream applications as well as their different evaluation methods. Finally, we identify the major challenges in developing SpeechLMs and outline potential directions for future research. We hope this survey will illuminate the field and assist the research community in creating more powerful Speech Language Models.